A firm's dividends have grown over the last several years. 3 years ago the firm paid a dividend of $2. Yesterday it paid a dividend of $7. What was the average annual growth rate of dividends for this firm?

Round the answer to two decimal places in percentage form.
Answer #1

The dividend paid 3 years back was $2.

The dividend paid after 3 years is $7.

So, the growth rate of dividend is :

= 2(1 + growth rate)^3 = 7

Growth rate = ($7/ $2) ^ 1/3
Growth rate = 51.82%

It can also be computed as ,

PV = ($2)

FV = $7

N = 3 YEARS

COMPUTE I/Y

So, I/Y = 51.83